# Which Similarity-Sensitive Entropy (S-entropy)?

### v3

@@ 3.1 Datasets and definitions
The following imaging and tabular datasets were used for benchmarking LCR and VS:
@@
  • Image datasets. 22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [36]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [33]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [21], the NIH ChestXRay dataset [34]; COCO (Common Objects in Context) [18]; and MIDRC COVIDx CXR-4 [35]. For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
@@
  • Tabular datasets. After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [11]. For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
@@
Definitions (Fig. 1).The terms “system” and “dataset” are used interchangeably. The elements of image and tabular datasets are single images and single rows, respectively. Each element appears only once in these datasets; i.e. each element is also a unique element.
3.2 Similarity matrices
@@
where $k$ is a parameter that was varied during experiments to test for effects on results. This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [17]. Imaging datasets were first embedded into a two-dimensional space using UMAP [20] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq. 7 equivalent to the root-mean-square distance (RMSD). The default $k$ for tabular datasets was $1$.
3.3 Entropy and S-entropy calculations
@@
The sentropy Python package [23] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1). The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2). VS was calculated using the vendi‑score package [10]. This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR. Similarity matrices were calculated on a GPU-enabled workstation equipped with multi‑core CPUs. LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
3.4 Clustering
@@
HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [19] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset. HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
3.5 Numerical experiments
Following a previous approach [8], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize. A negative value indicates a counterexample. To meet the assumptions of this conjecture, $Z$ was constructed as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$. To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative. We assumed a uniform distribution of unique elements. Thus, we parameterized $Z$ by a $r\times n$ matrix $x$, from which we defined
$Z=w^{\intercal}w;w_{ij}=e^{x_{ij}}(\displaystyle\sum_{k}e^{2x_{kj}})^{-1/2}$ (8)
We minimized the difference $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for each combination of $q\in\{-10$, 2.7-2.7, 2-2, 1-1, 0.3-0.3, 0, $0.3$, 11, $2.7$, $3.4$, 44, 55, $2\pi$, $7$, 88, $9.1$, 10}10\}, $n\in\{3,5,10,20\}$, $r\in\{2,\lfloor n/2\rfloor,n\}$ if $n>3$ and $r\in\{2,3\}$ for $n=3$ (to avoid $r=1$). We also varied the distribution from which the initial guess fed to scipy.optimize was drawn. In total, 4420 minimization were done. The full Python code for performing these minimizations can be found in section A.1
3.6 Speed tests
@@
Run times were calculated using the Python timeit package with 5-fold replications and measured on a 2024 M4 Apple Mac mini with 64GB unified memory.
4 Results
4.1 LCR and VS of major imaging and tabular ML datasets
@@
We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits. The results below are for a random 10,000-image subset, taken for computational convenience. Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
@@
Fig. 2a displays a UMAP embedding of the data, coloring each digit by its label. HDBSCAN detects six clusters, successfully grouping together digits that appear visually similar (Fig. 2b). LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct from each other (Fig. 2c). Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster. Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose. Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10). Mathematically, this indicates that this MNIST subset possesses the same entropy as a hypothetical dataset containing 12-13 completely dissimilar elements (i.e. with zero pairwise similarity). This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements. However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑ or “ur”‑images, each actual image in the dataset being a linear combination of these. Fig. 3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig. 3b, gray). Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits. VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues. The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig. 3b.
@@
Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets and smaller (and sometimes no) differences for tabular datasets.
@@ 4.2 Trends with LCR and VS
Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging and (b) tabular datasets. Each point represents a dataset. All entropies are effective-number forms at $q=1$. Red line = linear regression fit; gray line = 1:1 (off the scale to the left for (a), far right). Select datasets are labeled. (c) UMAPs for labeled imaging datasets. Gray points = unclustered elements.
@@
Figure 6: LCR vs. VS at $q=1$ at a very small half distance, an intermediate half-distance for which the correlation is low, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets. For tabular datasets, the intermediate half-distance is the one simultaneously minimizing the $R^{2}$, the correlation, and maximizing the covariance at $q=1$. For image datasets, the intermediate half-distance does not coincide with the global minimum of the correlation at $q=1$, because VSs and LCRs already come very close to 10,000 at that global minimum, so we chose a smaller value. (a) and (c): red line = linear regression fit; gray line = 1:1. Each point represents a dataset. (b) and (d): vertical gray lines = $k$s for (a) and (c).
@@
4.3 Bounds on LCR and VS
In Figs. 5 and 6, VS values were consistently higher than LCR, and both measures converge to 1 as the half‑distance decreases. It is therefore reasonable to ask whether the inequality $VS\geq LCR$ holds universally. We formalized this question as the following conjecture:
Conjecture 1.
Let $Z$ be an $n\times n$ ($n>0$) PSD similarity matrix, i.e. a PSD symmetric matrix with entries between 0 and $1$ whose diagonal entries equal $1$. Denote LCR and VS at order $q$ as $D_{q}(Z,p=\frac{1}{n})$ and $\text{VS}_{q}(Z)$, respectively, on the uniform distribution. Then for all $q\in[-\infty,+\infty]$
$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (9)
Note that in the context of this conjecture, since we are considering the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal. We prove this conjecture for $q=2,3,$ and $\infty$ in Appendix A.2. We consider the existence of proofs for these several $q$ to be strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not, for example, at $q=2.5,4$, or $27.3$. The following lemma supports this conjecture holding for all $q$:

### v4

@@ IV.1 Datasets and definitions
The following imaging and tabular datasets were used for numerical experiments comparing LCR and VS:
@@
  • Image datasets. 22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST [23], Fashion-MNIST [53], CIFAR-10 and CIFAR-100 [22] (from the torchvision [31] Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [54, 55]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [48]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [35], the NIH ChestXRay dataset [50]; COCO (Common Objects in Context) [29]; and MIDRC COVIDx CXR-4 [52]. For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset and the NIH ChestXRay dataset were downscaled to $250\times 250$ pixels, and images in the COCO dataset were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
@@
  • Tabular datasets. After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [16]. For each dataset, identifier columns and any features containing non-numeric strings or time-series data were removed.
@@
The terms “system” and “dataset” are used interchangeably (Fig. 1). The elements of image and tabular datasets are single images and single rows, respectively. Each element appears only once in these datasets; i.e. each element is also a unique element.
IV.2 Similarity matrices
@@
where $k$ is a parameter that was varied during experiments to test for effects on results. This similarity measure was chosen for its simplicity and for being provably positive-definite (PSD), guaranteeing valid similarity matrices for VS (see Section VI); LCR has no such restriction [27]. Imaging datasets were first embedded into a two- or three-dimensional space using uniform manifold approximation and projection (UMAP) [34] and the default $k$ taken to be $2^{-1/2}$ (varying random seeds to verify robustness), making the norm in Eq. 37 equivalent to the root-mean-square distance (RMSD). UMAP embeddings avoid the high dimensionality of raw pixel values, which has adverse effects on distance-based algorithms [2]. Note that UMAP projections are most often chosen to be in two dimensions. Empirically, UMAP embeddings generally place similar images near each other and different images further from each other (e.g. Figure 2), rationalizing the use of these embeddings in the literature and validating their use here. We present 2D UMAP results in the main text and 3D in the appendix A.
Unlike images, the dimensionality of the tabular datasets used here is low, rationalizing direct Euclidean distance. The default $k$ for tabular datasets was $1$.
IV.3 Entropy and sentropy calculations
@@
The sentropy Python package [37] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section II.1). The package accepts a user-provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon-type LCR) for direct comparison with the default VS (see Section II.2.2). VS was calculated using the vendi-score package [14, 39]. This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR. Similarity matrices were calculated on a GPU-enabled workstation equipped with multi-core CPUs. LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
IV.4 Clustering
@@
HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) [33] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset. HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well-suited for heterogeneous data.
IV.5 Speed tests
@@
As a practical comparison, empirical run times for sentropy and vendi-score were calculated using the Python timeit package with 5-fold replications and measured on a 2024 M4 Apple Mac Mini with 64GB unified memory.
IV.6 LCR and VS of imaging and tabular datasets
@@
We first compared LCR to VS, both at $q=1$, on a simple, well-known imaging dataset: MNIST handwritten digits. The results below are for a random 10,000-image subset, taken for computational convenience. Both quantities were expressed in their effective-number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state-of-the-art clustering algorithm HDBSCAN.
@@
Fig. 2a displays a UMAP embedding of the data, coloring each digit by its label. HDBSCAN detects six clusters, successfully grouping together digits that appear visually similar (Fig. 2b). LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct from each other (Fig. 2c). Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster. Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly when clusters are somewhat loose. Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10). Mathematically, this indicates that this MNIST subset possesses the same entropy as a hypothetical dataset containing 12-13 completely dissimilar elements (i.e. with zero pairwise similarity). This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the sentropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements. However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen- or “ur”-images, each actual image in the dataset being a linear combination of these. Fig. 3a displays the 100 eigen-images with the largest eigenvalues, which dominate the total entropy (Fig. 3b, gray). Some of these eigen-images resemble recognizable digits, whereas others resemble superpositions of multiple digits that do not otherwise correspond to anything obviously familiar. VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen-images at frequencies given by the eigenvalues. The total number of eigen-images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig. 3b.
@@
Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets (embedded in 2D UMAP projections) and smaller (and sometimes no) differences for tabular datasets (in their native space).
@@ IV.7 Trends with LCR and VS
Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging (2D UMAP; 10 different random seeds) and (b) tabular datasets. Each point represents a dataset. All entropies are effective-number forms at $q=1$. Red line = linear regression fit; gray line = 1:1 (off the scale to the left for (a), far right). Select datasets are labeled. (c) UMAPs for labeled imaging datasets, colored by cluster. Gray points = unclustered elements.
Figure 6: LCR vs. VS at a very small half-distance, an intermediate half-distance with low correlation, and large half-distance for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets. The scatterplots in (a) and (c) are for $q=1$. Red line = linear regression fit; gray line = 1:1. Each point represents a dataset. (b) and (d): vertical gray lines = the $k$ values plotted in (a) and (c).
@@
V Discussion
Quantifying the information contained in a system is essential for virtually any complex system. This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative. This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when the elements each appear only once, this distribution is flat (Fig. 1): there is no additional information there. Traditional entropy is used to measure the distribution of class sizes, which often vary; this is class balance. However, there is no obvious way to use class balance to measure the information contained within or between classes, which is a property of the elements themselves.
V.1 Sentropy as a measure of dataset composition
Conceptually, the main difference between clustering and sentropy is that sentropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0. The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership in that cluster and thereby contributing to the size of the cluster. This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster to which many unique elements belong fairly strongly would be considered “ordinary” (Fig. 2c). This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
How do LCR and VS compare with the number of clusters (as identified by HDBSCAN)? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs. 1, 5c). Indeed, many of the datasets we examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute to the HDBSCAN total).
In contrast, for imaging datasets, LCR and VS were typically 1-2 orders of magnitude lower than the number of clusters, indicating substantial similarity among elements across clusters, which hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of sentropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; a limitation of our study is that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity-sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search is unlikely to meaningfully affect the findings or conclusions presented.
V.2 The effect of scaling $Z$
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity- matrix values are scaled, which we illustrated by introducing the half-distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve in the middle of this range, while maintaining the anchoring at each end (at 0 and 1).
Why should one need to scale the similarity between elements? Should scaling be permitted at all? When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the external knowledge defines and thereby constrains the similarity function a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [5]. This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean-distance-based similarity measures used here (across different embeddings), there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$. For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc). Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context/are maximally informative about the system. Such choices arise often in measurement. For example, the parameter $q$ is an analogous dial in traditional entropy (as well as sentropy), allowing investigators to adjust the emphasis on frequency. As with $k$, prior information can constrain choice of $q$; for example, when the measurement has to do with collision frequency, $q=2$ should be used, consistent with $q=2$ corresponding to Simpson’s index. Whether specific values of $k$ might similarly be linked to known quantities of interest is a question left for future investigation.
A key feature of scaling is that it interpolates sentropy between 1 (for small $k$) and traditional entropy (for large $k$). When $k$ is small, the off-diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element. Conversely, for large $k$ the off-diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered. In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective-number form of traditional entropy equals the dataset size. Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$. Other scaling approaches may exhibit the same property.
Which $k$ should one choose? Qualitatively, the answer is a value roughly midway between the extremes (Fig. 6). However, our empirical studies indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for the tabular data sets and $3.4$ for the imaging data sets we tested—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$. Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section VI). The dependence on $k$, however, was quite strong.
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non-trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, from a quantitative perspective, the information in a system is in the eye of the beholder.
VI Conclusion
So, which sentropy: LCR or VS? Our results support defaulting to LCR when the goal is to measure the similarity-adjusted entropy of a system, and to prefer VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur-elements/eigen-elements, or (ii) when the system possesses a quantum-mechanical character (see below). If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non-trivial range (see Section V.2) to maximize their independence and thereby their combined informativeness.
We note that, quantitatively, VS $\geq$ LCR for any given $k$ and $q$. We proved this relationship for all $q\geq 0$ and for $q<0$ when $Z$ is full rank. This relationship can be rationalized by recognizing that, although VS is a similarity-sensitive entropy, its computation includes a traditional entropy component: after incorporating similarity to derive the eigen-elements, VS is the traditional entropy of the resulting eigenvalues; traditional entropy is always greater than or equal to similarity-sensitive entropy.
From a practical perspective, LCR has three advantages. First, its time complexity is $\mathcal{O}{(n^{2})}$ whereas computing eigenvalues—which is required to obtain the VS—has time complexity $\mathcal{O}{(n^{3})}$ [14]. Empirically, we observed a speed advantage of the sentropy package over the vendi-score package even when all elements were unique, as is often the case in ML datasets (Fig. 8). In addition, computing the LCR is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements.
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non-PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [26]). Thus, PSD-ness is irrelevant for LCR but must be checked to apply VS.
And third, LCR in its various forms [8, 9] (Table 1) is part of an extensively characterized framework [27] that provides similarity-sensitive analogs of mutual information, relative entropy, and cross-entropy and offers methods for partitioning the overall similarity-sensitive entropy ($\gamma$) into within-class ($\alpha$) and between-class ($\beta$) components. In entropic terms, $\alpha$ is the (effective-number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The sentropy package computes $\alpha$, $\beta$, and $\gamma$.)
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [38]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class-level similarity arises purely from element-level similarities within the classes. In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
In sum, LCR and VS provide complementary descriptions of similarity-sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay. We anticipate that the deeper understanding of their relationship and properties will aid such investigations.
VII Acknowledgements
Appendix A Comparison to 3D UMAP embedding
Fig. 9 shows results of the same analysis as in the right-hand panel of Fig. 5a but with the UMAP embeddings in three dimensions instead of two. This supports the lower-dimensional observations that VS and LCR tend to correlate to each other, but do contain independent information ($R^{2}=0.37$).
Figure 9: Results from 10 replicates with different random seeds for each dataset. Legend as in Fig. 5a.
Appendix B Other mathematical results
1 has an interesting consequence which is not used in our proof of 1, namely
Corollary 1.
Let Z, $\lambda_{i}$, and $r_{i}$ be as in 1. Denote by $r_{(i)}$ the $i$th row-sum of $Z$ ordered from largest to smallest, so that e.g. $r_{(1)}=\max_{i}r_{i}$. Then $r_{(i)}\geq\lambda_{i}$ for all $i$ in $\{1,\ldots,n\}$
Proof.
This is an immediate consequence of 1 and the Weyl monotonicity theorem [51]. ∎
This would seem a good basis for proving 1 except $VS_{q}$ depends on $\sum_{i}\lambda_{i}^{q}$, but $D_{q}\left(Z,\frac{1}{n}\right)$ depends on $\sum_{i}r_{i}^{q-1}$ rather than $\sum_{i}r_{i}^{q}$. It is clear, however, that 1 leads to some bound involving $VS_{q}$ and $D_{q+1}$. Further computation reveals that the relevant bound is
Corollary 2.
$\left[VS_{q}(Z)\right]^{1-q}\leq n^{\frac{1+q}{q}}[D_{q+1}]^{-q}$ (38)
for $q>0$;
$VS_{q}(Z)\begin{cases}\leq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}}\text{ for }0<q<1\\ \geq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}}\text{ for }q>1\end{cases}$ (39)
Proof.
This is a direct consequence of 1. Since $r_{(i)}\geq\lambda_{i}$, and since $x\rightarrow x^{q}$ is monotonically increasing for $q>0$, we have for each $i$ that
$r_{(i)}^{q}\geq\lambda_{i}^{q}$ (40)
$\Rightarrow$ (41)
$\sum_{i}\lambda_{i}^{q}\leq\sum_{i}r_{i}^{q}$ (42)
However, for $q\neq 1$,
$\sum_{i}r_{i}^{q}$ $=n^{1+q}[D_{q+1}]^{-q}$ (43)
$\sum_{i}\lambda_{i}^{q}$ $=n^{q}\left[VS_{q}(Z)\right]^{1-q}$ (44)
thereby establishing the claim for $q\neq 1$. The $q=1$ case then follows by continuity in $q$. ∎
Note that this bound is weaker than 3 for $q>1$, since $r_{i}^{q-1}\leq r_{i}^{q}$. On the other hand, for $0<q<1$ it provides an upper (rather than lower) bound on $VS_{q}(Z)$, since 1q>01-q>0, but this bound is also weaker than this simple bound that $VS_{q}(Z)<n$. As an illustrative example, consider the case of $q=\frac{1}{2}$. Then 2 leads to
$VS_{\frac{1}{2}}\leq\frac{n^{6}}{D_{\frac{3}{2}}\left(Z,\frac{1}{n}\right)}.$ (45)
This is clearly much weaker than $VS_{\frac{1}{2}}<n$, since $D_{\frac{3}{2}}\left(Z,\frac{1}{n}\right)\leq n$ means the right hand side is never smaller than $n^{5}$. In fact, this is in a sense the best case. Taking the inequality at general $0<q<1$, namely
$VS_{q}(Z)\leq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}},$ (46)
$D_{q+1}\leq n$ means the right hand side is bounded below by $n^{\frac{1+q-q^{2}}{q(1-q)}}$. Minimizing the exponent (with the restriction that $0<q<1$) gives $q=\frac{1}{2}$, so this right hand side is always bounded below by $n^{5}$, where as the left hand side is bounded above by nn—a very weak bound.
Another bound can be proven: